Replace ☐ in each of the following by the correct number:
 2 / 7 = 8 / ☐

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given
2 / 7 = 8/☐
2 ×☐=7 ×8
☐= (7 × 8) / 2
= 28

Thanks